A flag of mass 2.9 kg is supported by a single rop
Answer #1

T- tension; A -angle w horizontal

Assuming flag mass is vertical,

vertical component T sin A = 2.9 * g ----(1)

horizontally component T cos A = 14 ----(2)

deviding (1) /(2)

tan A = 2.9*g / 15 = 2.9 * 9.81/15

tan A = 1.8966

A = 62.2 deg

-------------------

Substituting A into (2)

T cos 62.2 = 15

T = 32.16 N